A Complete Bibliography of Publications in Advances and Applications in Statistical

Sciences

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

20 March 2019 Version 1.00

Title word cross-reference

2 [LLF10]. 3/2 [LLF10]. AR(1) [CN11]. C^{∗} [Cho18]. M [WW11]. P [HGXA10, Cho18]. ρ [RL10]. S = 1 [LLF10]. SU(2) [SP11]. t [Wan11].

-adic [Cho18]. -algebras [Cho18]. -distributions [Wan11]. -estimation [WW11]. -value [HGXA10].

accelerated [ZZ10]. accurate [RL10]. actions [Mar10b]. activity [NBC11]. adaptive [Hus10]. adic [Cho18]. admission [Hur11]. affecting [NFZ11]. after [CN11]. Agent [BPR^{+}10, CW10]. agent-based [CW10]. aggregation [LM11a]. agricultural [BPR^{+}10]. AHP [SM11]. air [Mor10]. algebras [Cho18]. algorithm [AET10, Fok12]. allocation [SMLS14]. allocative [NG11]. almost [YW10]. among [LB10]. analyses [HGXA10, Hus10]. analysis [AA16, Ing10, LB10, Luo12, MAAH13, PS11a, PVF11, RA10, SC10, TBLW10, WSNP10, WXZ10]. analytic [Cho18]. analytical [CGGS11]. analyzing [Fot10, GGJ10, Mas11]. annuity [PS11a]. anticipating [Yin11]. Application [KP10a, AN13, AGMR10, Baš11, EAM17, GG13, LB10, NGGJ10, Wan09, WSNP10, WN10b]. applications [Wan10, WL10]. approach [Ber11, GGJ10, Mor10, RRA11, SP10]. Approximate [Thi12]. approximation [AA16]. approximations [FS10a]. arbitrage [Gio11, WXZ10]. Argentine [BPR^{+}10]. assessing [NBC11]. assessment [BBN^{+}10]. athletes [Hur11]. attendance [HDS10]. autoregressive [LM11a]. Axiom [NN10]. axioms [NN10].

backward [LA10]. based [BPR^{+}10, CW10, He12, KMV10, Mas11, RPGM10, Vas10, Wan10, WW11, YKT12]. Bayesian [FG10, Hus10, KP10b, NBC11, RRA11, ZZ10]. bears [Par10]. Behavior [Bou10, SS10]. benefit [WN10a]. Benefits [LM15]. Bernoulli [FS10a]. Beta [Fok12, FC10]. between [FC10, Gio11]. beyond [Mac11]. binary [DHK09, Fot10]. bioassay [Che10]. bivariate [AT17, LPPY10]. block [FFNM12]. bolts [Mac11]. Bootstrapped [Hus10]. Box [AA16, WW11]. Brazil [PVF11]. breast [LAB^{+}10]. Brownian [XK11]. Bulls [Par10]. Business [LM15].

Calculation [Hut09]. Can [NN10]. Canada [Hur11, TBLW10]. cancer [LAB^{+}10, MCB10]. care [LM11b]. Carlo [AN12, SM15]. case [HW10, Hur11, LM15, LTC10, PVF11, Shi10, Spa11a, Spa11b]. censored [KP10a, Thi10, TC10]. censoring [BG11]. Central [Gri14, Gri16, YW10]. chain [PLFRKT11]. chains [AET10]. changes [BPR^{+}10, NBC11, WW11]. characterization [FZSR10]. Chaudhuri [Bou10]. child [WSNP10]. childhood [MCB10]. Choice [NN10]. choices [Mar10b]. churn [FG10]. cities [MLMB10]. claim [Thi12]. class [Kim11, LB10, WSNP10]. Classification [Tar10, Luo11]. client [Hur11]. client-statistician [Hur11]. clinical [Hus10]. Cluster [SC10]. clustered [GGJ10]. clustering [SMLS14]. cocaine [Wan10]. coefficient [RL10]. cointegrated [LM11a]. collective [Ber11, SS10]. College [Hur11]. Combining [TWW10]. common [Sel12]. commonly [DHK09]. commutative [FFNM12]. comparative [DHK09]. compared [AN12]. comparing [Mar10c]. comparison [DKH^{+}12, SM15]. competing [BG11]. complex [OPA11]. compounds [LAB^{+}10]. computing [AET10]. concepts [PF10]. conditional [Spa11b]. conditioned [AET10]. conditions [CCGF10, ZRQ11]. considerations [BD11]. consistency [CCGF10, ZRQ11]. consistent [AN12, HW10]. Constructing [CGGS11]. Consumption [GG13]. containing [Sel12]. contingency [TTT09, YKT12]. continuous [Mar10b]. contributions [WN10b]. control [He12, Pen13, RA10]. convergence [Gri16, Sto10]. Convolution [FS10a]. Copenhagen [LM15]. copulas [EAM17]. corpuscle [HW10]. correlated [DHK09]. correlation [RL10]. count [Sel12]. counterflow [Par10]. course [LM15]. Cox [WW11]. crack [Wan10]. crack-cocaine [Wan10]. criterion [Mas11]. criticisms [Spa11a]. crossing [FFNM12]. cubic [FC10]. current [TWW10]. curvature [BFSY10]. curve [BFSY10]. cycle [GG13].

daily [WW11]. data [Dan10b, DHK09, EAM17, Fot10, GGJ10, Gio11, HGXA10, KP10a, KMV10, OPA11, OR10, RRA11, SMLS14, Sel12, SM15, TBLW10, Thi10, TC10, TWW10, WN10b, WW11]. dating [ZA10, ZA11]. DEA [Shi10]. Dealing [WN10a]. Decision [MSV10, SP10, He12, SP11, ZA11]. decisions [Hur11]. decomposition [SM11]. delta [TTT09]. demand [Kur11]. density [XK11]. dependence [LPPY10]. Derivative [MPT10]. derived [Dra11]. describing [Sel12]. desertion [KMV10]. designs [Mas11]. detection [LAB^{+}10, LZ11, WW11]. determinants [Hut09]. diabetes [Luo12]. diagrams [RL10]. differences [Sel12]. differential [LZ11]. digraphs [BHS11]. dimensional [HGXA10]. directed [LLF10]. Dirichlet [Kar09]. disability [LM11b]. discovering [RPGM10]. discovery [KMV10]. discrete [Dra11, Mar10b]. dispersion [Sel12]. Dissipative [SP11]. Distance [FC10, YKT12]. distribution [Dra11, Fot11, HGXA10, KP10c, Sel12, Thi10, Thi12, Var10]. distributions [AET10, Bak10, HA10, Wan11]. divergence [YKT12]. down [Lef10]. drug [LB10]. during [GG13]. dynamic [PVF11]. Dynamics [Pro10, PS11b].

econometric [Kur11]. Economic [Pro10, PS11b]. education [Mac11]. Effect [AA16, CL11, Fot11]. effects [Fot11]. efficacy [HDS10]. efficiencies [JR10, NG11]. elasticity [CW10]. election [PVF11]. elements [Cho18]. elementwise [ZRQ11]. elite [Hur11]. emission [Baš11]. Empirical [BBN^{+}10, PS11a, Var10]. epidemics [MLMB10]. equality [PT11, TT09]. equation [LA10]. equations [CGGS11, GGJ10]. equivalence [Pen13]. error [AGMR10, Hut09, NG11, ZRQ11]. Estimate [Che10, RL10]. estimating [GGJ10]. Estimation [FFNM12, NG11, SM11, AN12, BG11, Fot11, HW10, Maz10, OR10, RRA11, Thi10, WW11, ZZ10]. estimator [ZRQ11]. estimators [CCGF10, FS10b]. Evaluating [HDS10]. evaluation [MPT10]. even [FS10b]. event [Fot10, TC10]. exam [LM15]. examination [AN12]. example [Spa13]. exchange [PVF11]. exhaled [LAB^{+}10]. Expected [Maz10]. expenditure [GG13]. experiment [Spa13]. Explicit [LA10]. Exploring [CL11]. exponential [Bak10, HA10, KP10c]. exponentiated [Bak10]. Export [AN13]. Export-led [AN13]. expression [SMLS14]. extension [Dan10a]. extreme [Thi10].

factor [WSNP10]. factors [Luo12]. failure [ZZ10]. fallacious [Spa13]. Farima [AGMR10]. fat [NG11]. fat-tailed [NG11]. fees [BD11]. films [Wal10]. finance [Maz10]. financial [AA16, Gio11]. First [XK11, LM11a]. fit [EAM17]. fixed [He12]. fixed-period [He12]. flight [CL11]. forecasting [Mor10]. Foreword [Pro10, PS11b]. formula [Yin11]. forward [Yin11]. Fractal [FS11]. Fraenkel [NN10]. frequentist [Spa13]. frictional [WXZ10]. function [Var10]. functional [OR10, Var10]. functions [SM15]. fund [BD11]. funds [PVF11]. future [KP10b]. fuzzy [MSV10, RRA11].

gamma [Bak10, BBN^{+}10]. Garch [AGMR10]. Gauging [Wan09]. GDP [Baš11]. gene [EAM17, LZ11, SMLS14]. General [Wal10, Yin11]. Generalized [FS10b, Wan11, BBN^{+}10, Dan10a, GGJ10, HA10, WXZ10]. generation [DHK09]. geometric [XK11]. geometry [Mar10a]. GLM [Dan10a]. Goodness [EAM17]. Goodness-of-fit [EAM17]. ground [Mor10]. group [Wan10]. group-based [Wan10]. Growth [PF10, AN13, MR11, Wan09, Wan10].

Hamiltonian [SP11]. hazard [LTC10]. health [BBN^{+}10]. healthcare [TBLW10, WN10a]. help [CW10]. heterogeneity [WSNP10]. heteroscedasticity [Kar09]. heteroskedasticity [AN12]. Hierarchical [NBC11]. high [HGXA10, LB10]. high-risk [LB10]. historical [TWW10]. history [Fot10]. Hollywood [Wal10]. homogeneity [YKT12]. household [GG13]. HRMs [He12]. hybrid [LTC10]. Hydrophilicity [Kim11]. hyperinflation [SS10]. hyperstructures [MSV10]. hypotheses [Spa11a]. hypothesis [AN13, Hut09].

ideological [AVD11]. II [KP10a]. ill [AET10]. ill-conditioned [AET10]. Illustrations [HGXA10]. implementation [PF10]. improper [Spa13]. incidence [Dan10a, MCB10]. income [GG13]. incomplete [SM15]. increments [Var10]. independence [AT17]. independent [FS10a]. index [Kar09]. indices [FS11]. individual [Ber11]. induced [Cho18, Fok12]. inefficiency [FZSR10]. inference [KP10b, Spa11b]. inflation [Kur11]. inflation-demand [Kur11]. Information [OPA11, Mas11, Vas10]. insensitive [Bou10]. instructions [MAAH13]. insurance [LM11b]. integral [CGGS11, Yin11]. Integration [Shi10]. intensity [Baš11]. interactions [Ber11]. interface [Hur11]. Internet [Lef10]. interval [Gio11, TC10]. interval-censored [TC10]. intervals [Gio11]. invariant [Mar10a]. inventory [He12]. Investigating [WSNP10]. involving [KP10c, Shi10]. Iran [GG13, MR11]. Ising [LLF10]. issue [Pro10, PS11b, WL10]. Issues [RPGM10]. Itô [Yin11]. Italian [RA10]. IV [Pro10].

Japan [Kur11]. Jenkins [AA16]. joint [Fot10, KP10c].

knowledge [RPGM10]. Kolmogorov [LA10].

large [Sto10]. latent [Ing10, LB10, Wan09, WSNP10]. law [Sto10, Var10]. LCA [NGGJ10]. LDA [NFZ11]. leaders [FZSR10]. least [AN12, AN13, ZRQ11]. led [AN13]. lessons [Mac11]. level [Mor10]. levels [Sel12]. Lévy [Yin11]. life [GG13, KP10a, PS11a]. Lifetime [TBLW10]. like [Cho18]. likelihood [CCGF10, TC10, TWW10]. limit [Gri14, Gri16, Var10, YW10]. limitation [NBC11]. limiting [AET10]. line [BHS11]. Linear [Baš11, WN10b, BBN^{+}10, Dan10a, SP11]. link [PLFRKT11]. Local [TC10]. long [LM11b]. Longevity [PS11a, LM11b]. Longitudinal [Wan10, Dan10b, Fot10]. Lotka [AVD11].

macroeconomics [LM15]. making [MSV10, SP10, SP11]. malaria [Dan10a]. management [BD11, Vas10]. managerial [Shi10]. Managing [Hur11, LM11b]. mandatory [HDS10]. manufacturing [PLFRKT11]. Mapping [MCB10]. marginal [TTT10, YKT12]. market [FZSR10, FS11, RA10, ZA10, ZA11]. markets [Gio11, WXZ10]. Markov [AET10, He12, XK11]. martingales [Gri14]. mathematical [Ber11]. matrix [CGGS11]. maximal [Mar10a]. maximum [CCGF10]. mean [WW11]. means [Mar10c, PT11]. Measure [YKT12, Mar09]. measurement [Dan10a, Mac11, ZRQ11]. measures [MSV10]. median [FS10b]. medical [Mac11, WN10b]. meet [NN10]. Meeting [Pro10, PS11b]. method [AN12, FZSR10, Luo11, ZZ10]. methodology [AA16]. Methods [WL10, DHK09, DKH^{+}12, LZ11]. microarray [LZ11]. middle [Mar10b, NGGJ10]. mild [ZRQ11]. Military [Hur11]. Mining [LAB^{+}10, KMV10]. Misplaced [Spa11a]. mission [Vas10]. misspecification [CL11]. mixed [BBN^{+}10, Spa13, WN10b]. mixing [Fot11]. mixture [Bak10, PF10, WL10]. model [AVD11, Baš11, BG11, Bou10, CN11, Dan10a, Dra11, DKH^{+}12, FLP13, FS11, Fot10, Ing10, Kar09, KP10a, Kur11, LLF10, LTC10, Mar10b, PLFRKT11, SP10, Spa11b, Spa13, TTT09, ZZ10, ZRQ11]. Modeling [Kar09, MLMB10, Vas10, CW10, OPA11, PF10, Shi10, Wan09]. Modelling [LM11b]. models [BBN^{+}10, FFNM12, FG10, Fot11, LM11a, NBC11, NG11, Wal10, Wan10, WL10, WN10b]. Monte [AN12, SM15]. most [DHK09]. motion [XK11]. Moving [Mac11]. MSE [JR10]. multi [PLFRKT11, TTT10]. multi-way [TTT10]. multicollinearity [WN10a]. Multifractal [FZSR10]. multilevel [NGGJ10]. multinomial [TT09]. multivariate [FLP13, Mar09, Mar10c, Var10, ZRQ11].

N [Spa11a, Kim11]. n-tio [Kim11]. Necessary [CCGF10]. need [MAAH13]. Network [RA10]. networks [LLF10]. Newey [AN12]. nexus [Kur11]. Neyman [Spa11a]. Nigeria [AN13]. no [PT11, WXZ10]. no-arbitrage [WXZ10]. non [SP11]. non-linear [SP11]. nonlinear [Mor10]. Nonparametric [BG11, Mor10]. nonstandard [Var10]. northern [MCB10]. note [Dra11, KP10c, Mar09, YW10, ZRQ11]. numbers [Sto10]. nuts [Mac11].

occluded [HW10]. oil [FLP13]. one [BHS11, PT11]. one-unit [BHS11]. operator [OR10]. opinions [Mar10b]. Optimal [He12]. optimization [Luo11]. option [Mar10b]. options [WN10a]. oral [BBN^{+}10]. order [Gio11, LM11a]. ordered [TT09]. ordering [Pen13]. ordinary [AN12]. organic [LAB^{+}10]. organization [PLFRKT11]. orthogonal [FFNM12]. outcomes [BBN^{+}10, Wan09]. ownership [RA10].

P [Spa11a]. Pampas [BPR^{+}10]. panel [Dan10b]. parabolic [Che10]. parallel [KP10b, Wan09]. parameters [Fot11]. parametric [DKH^{+}12]. parsimonious [Kur11]. partitioning [AET10]. passage [XK11]. Pattern [KMV10]. patterns [Ber11]. Peano [NN10]. Pearson [Spa11a, RL10]. pedestrian [Par10]. peer [NGGJ10]. penalized [Luo11]. pension [BD11]. People [SS10]. perceptions [Ing10]. performance [NFZ11]. period [He12, PLFRKT11]. Petrov [Sto10]. photocatalytic [Kim11]. plan [PLFRKT11]. point [RRA11, TTT10]. point-symmetry [TTT10]. points [FC10]. Poissonian [KP10c]. policy [He12]. pollutants [Mor10]. population [MR11]. Portfolio [BD11]. Portugal [MCB10]. positive [LPPY10]. possible [NBC11]. posteriori [Maz10]. potency [Che10]. power [YKT12]. power-divergence [YKT12]. precipitation [WW11]. prediction [CL11]. predictive [SMLS14]. Predictor [CN11]. Preface [WL10]. preferences [Shi10]. preliminary [CN11]. preparation [LM15]. prevalence [NBC11]. prices [FLP13]. Pricing [Gio11]. priorities [SM11]. private [LM15]. Probabilistic [MAAH13]. probabilities [BHS11, Thi12]. probability [Dra11]. problem [HW10]. procedure [Bou10]. Proceedings [Pro10, PS11b]. process [AGMR10, He12, Kar09, RPGM10]. processes [Yin11]. production [PLFRKT11]. products [FLP13]. profile [Ing10]. projection [MR11]. propagation [MLMB10]. properties [AN12]. property [Wan11]. proportions [TT09]. puzzle [CW10].

Quantile [Thi10]. Quantum [ZA10, ZA11, SP10]. quick [RL10].

random [FC10, Fot11, FS10a]. randomized [Bou10]. ranked [Bou10]. Rasch [Dra11, DKH^{+}12]. Rate [Sto10, Hut09, MR11, PVF11, Wal10]. rather [RL10]. Rayleigh [KP10a]. reasoning [Spa13]. recombining [MPT10]. recurrent [CGGS11, TC10]. refined [FLP13]. region [MCB10]. regional [SC10]. regression [AGMR10, Baš11, BFSY10, BG11, JR10, OR10]. regressional [Wan11]. regressions [Mar10a]. related [Wan09]. relation [Luo12]. relationships [WSNP10]. relative [Che10]. reliability [Mac11, Vas10]. remission [KP10a]. repairer [BHS11]. repeated [Dan10a]. replication [PT11]. representation [LPPY10]. research [BBN^{+}10]. response [Bou10]. responses [KP10b]. restrictions [Shi10]. return [Wal10]. Revisiting [Spa11b, Spa13]. ridge [JR10]. right [BG11]. risk [LB10, PS11a]. risks [BG11, LM11b]. root [CN11]. Royal [Hur11]. ruin [Thi12]. rule [RPGM10].

safety [Ing10]. sample [AN12, FS10b, Mar09, PT11]. sampler [SMLS14]. sampling [Bou10]. scarce [OPA11]. scatter [RL10]. schemes [WN10b]. school [Ing10, NGGJ10, LM15]. science [SC10]. score [Mac11]. search [ZA10]. seasonal [CN11]. selection [BD11, Bou10]. Semicircular [Cho18].

Semicircular-like [Cho18]. Semiparametric [AGMR10, ZZ10]. semiquantum [SP11]. series [AA16, Mar10a, Tar10, WW11]. set [Bou10, KP10a, NN10, Vas10]. several [Pen13]. shifts [WW11]. short [Tar10]. simple [CN11, Pen13, Spa11a]. simulation [BPR^{+}10, Fot10, SM15]. simulator [FLP13]. simultaneously [Wan09]. single [Kar09]. single-index [Kar09]. size [Thi12]. sizes [FS10b]. skewness [Mar09]. slowing [Lef10]. small [AN12, FS10b]. Social [Pro10, PS11b]. society [OPA11]. software [Vas10]. solutions [CGGS11, LA10]. Some [Mar10a, BD11, EAM17, FS10a, Mac11, SM15]. space [FLP13]. spaces [FC10]. sparsity [Fok12]. Special [Pro10, WL10, PS11b]. specific [Vas10]. spheres [HW10]. spot [FLP13]. square [AN12, AN13, TTT09, YKT12]. squares [ZRQ11]. stable [Gri16, Wal10]. stage [AN13]. standby [BHS11, MAAH13]. state [BHS11, FLP13]. Statistical [LZ11, Luo12, KP10b, Spa13]. statistician [Hur11]. Steady [BHS11]. Steady-state [BHS11]. Steck [Hut09]. Stein [OR10]. Stenger [Bou10]. stochastic [MR11, MPT10, PLFRKT11]. stock [FZSR10, FS11, RA10]. stress [LTC10, SP10]. strong [Sto10]. Strongly [HW10]. structure [FFNM12, RA10]. structures [Cho18]. struggles [AVD11]. students [KMV10]. study [Bak10, DHK09, Fot10, NGGJ10, Shi10, Wan10]. style [PVF11]. sufficient [CCGF10]. sum [KP10c]. sums [Hut09]. supersaturated [Mas11]. supply [PLFRKT11]. sure [YW10]. survey [GG13]. survival [SM15, TWW10]. switching [XK11]. symmetrical [AT17]. symmetry [TTT09, TTT10]. system [BHS11, MAAH13, OPA11, Par10]. systems [BPR^{+}10, KP10b, Vas10, Pro10, PS11b].

tables [TTT09, TTT10, YKT12]. tailed [NG11]. taking [LM15]. taxicab [SM11]. teacher [WSNP10]. teacher-child [WSNP10]. technical [NG11]. technique [AN13]. Temporal [LM11a]. term [LM11b]. test [AT17, TTT10]. Testing [PT11, Pen13, TT09, DKH^{+}12, Spa11a]. tests [CN11, EAM17, Mar10c]. theorem [Gri16, YW10]. theorems [Gri14]. theory [AGMR10, Mac11, NN10, ZA11]. threshold [LTC10]. threshold-stress [LTC10]. time [AA16, Mar10a, TBLW10, Tar10, ZZ10]. times [KP10a, XK11]. tio [Kim11]. tools [Vas10]. total [ZRQ11]. Tracing [BFSY10]. Tracking [JR10]. trajectories [Wan09]. trajectory [Wan10]. transformed [WW11]. transition [LB10]. Transitions [LB10]. treatments [Pen13]. tree [Pen13]. trees [MPT10]. trend [WW11]. trials [Hus10]. trivariate [FC10]. tutorial [LM15]. two [AN13, BHS11, FC10, Hus10, MAAH13, Spa11a, TT09, Wan09]. two-repairer [BHS11]. two-stage [AN13]. type [KP10a, OR10].

uncertainty [CGGS11]. undocumented [WW11]. unemployment [GGJ10]. unequal [Bou10]. unified [LPPY10]. uniform [Spa11b, Var10]. unit [BHS11, CN11]. universality [Kim11]. university [KMV10]. unobserved [WSNP10]. use [LB10, Wan10]. used [DHK09]. Using [Ing10, AA16, BHS11, FLP13, FZSR10, HGXA10, MR11, MPT10, TWW10].

V [PS11b]. valuable [Mac11]. value [HGXA10, Thi10]. variables [Bou10, FS10a, KP10c, NFZ11]. varsity [Hur11]. vector [LM11a]. Verhulst [AVD11]. via [Hut09]. victimization [NGGJ10]. virtual [MLMB10]. VLV [AVD11]. volatile [LAB^{+}10]. volatility [MPT10]. Volterra [AVD11].

waiting [TBLW10]. warm [MAAH13]. wavelet [FZSR10]. way [TTT10]. weak [ZRQ11]. Weibull [LTC10, Thi12]. weight [CL11, Shi10]. weighted [GGJ10, TWW10, ZRQ11]. Welch [Spa11b]. West [AN12]. Wicksell [HW10]. without [WW11]. women [LB10]. would [CW10].

Zermelo [NN10].

References

Adiele:2016:EAF

[AA16] David F. Adiele and Vivian O. Adieles. Effect of approximation on financial time series analysis using Box–Jenkins methodology. Advances and Applications in Statistical Sciences, 11(1): 1–8, February 2016. CODEN ???? ISSN 0974-6811.

Aji:2010:PAC

[AET10] C. Affane Aji and H. Elalaoui-Talibi. A partitioning algorithm for computing limiting distributions of ill-conditioned Markov chains. Advances and Applications in Statistical Sciences, 1(2):511–519, February 2010. CODEN ???? ISSN 0974-6811.

Aneiros:2010:SRF

[AGMR10] Germán Aneiros, Wenceslao González-Manteiga, and Juan C. Reboredo. Semiparametric regression with a Farima–Garch error process: theory and application. Advances and Applications in Statistical Sciences, 3(1):83–112, September 2010. CODEN ???? ISSN 0974-6811.

Adiele:2012:MCE

[AN12] David F. Adiele and Joy C. Nwabueze. Monte Carlo examination of small sample properties of Newey West heteroskedasticity consistent estimation method compared to ordinary least square estimation method. Advances and Applications in Statistical Sciences, 8(1):1–10, November 2012. CODEN ???? ISSN 0974-6811.

Adiele:2013:ELG

[AN13] D. F. Adiele and J. C. Nwabueze. Export-led growth hypothesis in Nigeria: an application of two-stage least square technique. Advances and Applications in Statistical Sciences, 8(4):111–??, July 2013. CODEN ???? ISSN 0974-6811.

Amerise:2017:NST

[AT17] Ilaria L. Amerise and Agostino Tarsitano. A new symmetrical test of bivariate independence. Advances and Applications in Statistical Sciences, 12(2):31–53, May 2017. CODEN ???? ISSN 0974-6811.

Ausloos:2011:VLV

[AVD11] Marcel R. Ausloos, Nikolay K. Vitanov, and Zlatinka I. Dimitrova. Verhulst–Lotka–Volterra (VLV) model of ideological struggles. Advances and Applications in Statistical Sciences, 6(6):497– 503, 2011. CODEN ???? ISSN 0974-6811.

Bakoban:2010:SME

[Bak10] Rana Ali Bakoban. A study on mixture of exponential and exponentiated gamma distributions. Advances and Applications in Statistical Sciences, 2(1):101–127, 2010. CODEN ???? ISSN 0974-6811.

Basic:2011:LRM

[Baš11] Aleksandra Murks Bašič. Linear regression model: an application to GDP emission intensity. Advances and Applications in Statistical Sciences, 6(7):615–642, October 2011. CODEN ???? ISSN 0974-6811.

Baldi:2010:EAG

[BBN^{+}10] Ileana Baldi, Paola Berchialla, Vincenzo Notaro, Sandro Barone-Monfrin, Francesco Bassi, and Dario Gregori. Empirical assessment of gamma generalized linear mixed models for outcomes in oral health research. Advances and Applications in Statistical Sciences, 3(1):43–64, September 2010. CODEN ???? ISSN 0974-6811.

Baione:2011:PSM

[BD11] Fabio Baione and Paolo De Angelis. Portfolio selection and management fees: some considerations for a pension fund management. Advances and Applications in Statistical Sciences, 6(6): 373–391, September 2011. CODEN ???? ISSN 0974-6811.

Bertotti:2011:IIC

[Ber11] Maria Letizia Bertotti. From individual interactions to collective patterns. A mathematical approach. Advances and Applications in Statistical Sciences, 6(6):461–478, September 2011. CODEN ???? ISSN 0974-6811.

BenHur:2010:TCR

[BFSY10] D. Ben Hur, L. Frantskevich, E. Schechtman, and S. Yitzhaki. Tracing the curvature of a regression curve. Advances and Applications in Statistical Sciences, 4(1):31–56, November 2010. CO-DEN ???? ISSN 0974-6811.

Bordes:2011:NER

[BG11] Laurent Bordes and Kossi Essona Gneyou. Nonparametric estimation of regression in competing risks model with right censoring. Advances and Applications in Statistical Sciences, 6(1): 39–57, March 2011. CODEN ???? ISSN 0974-6811.

Bieth:2011:SSP

[BHS11] Bruno Bieth, Liang Hong, and Jyotirmoy Sarkar. Steady-state probabilities of a one-unit two-repairer standby system using line digraphs. Advances and Applications in Statistical Sciences, 5(2): 113–127, February 2011. CODEN ???? ISSN 0974-6811.

Bouza:2010:BRR

[Bou10] Carlos N. Bouza. Behavior of a randomized response procedure under unequal selection model of Chaudhuri–Stenger for insensitive variables under ranked set sampling. Advances and Applications in Statistical Sciences, 4(2):135–144, December 2010. CODEN ???? ISSN 0974-6811.

Bert:2010:ABS

[BPR^{+}10] F. Bert, G. Podestá, S. Rovere, A. Menéndez, X. González, F. Ruiz Toranzo, M. Torrent, M. North, C. Macal, P. Sydelko, E. Weber, and D. Letson. Agent based simulation of recent changes in agricultural systems of the Argentine Pampas. Advances and Applications in Statistical Sciences, 2(2):213–231, August 2010. CODEN ???? ISSN 0974-6811.

Cavazos-Cadena:2010:NSC

[CCGF10] Rolando Cavazos-Cadena and Graciela M. González-Farías. Necessary and sufficient conditions for the consistency of maximum likelihood estimators. Advances and Applications in Statistical Sciences, 3(1):13–42, 2010. CODEN ???? ISSN 0974-6811.

Chebrakov:2011:CAS

[CGGS11] Y. V. Chebrakov, C. B. García, J. Garcia, and V. V. Shmagin. Constructing analytical solutions with uncertainty for recurrent, matrix and integral equations. Advances and Applications in Statistical Sciences, 7(1):1–21, November 2011. CODEN ???? ISSN 0974-6811.

Chen:2010:ERP

[Che10] D. G. Chen. Estimate the relative potency in parabolic bioassay. Advances and Applications in Statistical Sciences, 2(1):1–17, May 2010. CODEN ???? ISSN 0974-6811.

Cho:2018:SLS

[Cho18] Ilwoo Cho. Semicircular-like and semicircular elements induced by p-adic analytic structures and C^{∗}-algebras. Advances and Applications in Statistical Sciences, 13(1):1–51, November 2018. CO-DEN ???? ISSN 0974-6811.

Chang:2011:EEW

[CL11] Jing Chang and Herbert K. H. Lee. Exploring the effect of weight misspecification on flight prediction. Advances and Applications in Statistical Sciences, 6(1):27–38, 2011. CODEN ???? ISSN 0974-6811.

Chiangpradit:2011:PSS

[CN11] Monchaya Chiangpradit and Sa-Aat Niwitpong. Predictor of simple seasonal AR(1) model after preliminary unit root tests. Advances and Applications in Statistical Sciences, 6(3):175–185, May 2011. CODEN ???? ISSN 0974-6811.

Chen:2010:EPW

[CW10] Shu-Heng Chen and Shu G. Wang. On the elasticity puzzle: would the agent-based modeling help? Advances and Applications in Statistical Sciences, 2(2):375–391, August 2010. CODEN ???? ISSN 0974-6811.

Dansu:2010:EGL

[Dan10a] Bolanle M. Dansu. An extension of generalized linear model (GLM) in repeated measurement of malaria incidence. Advances and Applications in Statistical Sciences, 2(1):69–88, 2010. CO-DEN ???? ISSN 0974-6811.

Dansu:2010:LPD

[Dan10b] Bolanle M. Dansu. On longitudinal and panel data. Advances and Applications in Statistical Sciences, 4(1):73–78, 2010. CODEN ???? ISSN 0974-6811.

Demirtas:2009:CSM

[DHK09] Hakan Demirtas, Donald Hedeker, and Kush Kapur. A comparative study on most commonly used correlated binary data generation methods. Advances and Applications in Statistical Sciences, 1(1):45–55, November 2009. CODEN ???? ISSN 0974-6811.

Draxler:2012:CPM

[DKH^{+}12] Clemens Draxler, Helmut Küchenhoff, Moritz Heene, Markus Bühner, and Julia Mang. A comparison of parametric methods testing the Rasch model. Advances and Applications in Statistical Sciences, 8(1):11–27, 2012. CODEN ???? ISSN 0974-6811.

Draxler:2011:NDP

[Dra11] Clemens Draxler. A note on a discrete probability distribution derived from the Rasch model. Advances and Applications in Statistical Sciences, 6(7):665–673, October 2011. CODEN ???? ISSN 0974-6811.

Elaal:2017:GFT

[EAM17] Mervat K. Abd Elaal, Samia A. Adham, and Heba M. Malaka. Goodness-of-fit tests for some copulas an application for gene data. Advances and Applications in Statistical Sciences, 12(1): 1–30, February 2017. CODEN ???? ISSN 0974-6811.

Fotouhi:2010:DBT

[FC10] Ali Reza Fotouhi and David P. Chu. Distance between trivariate beta random points in two cubic spaces. Advances and Applications in Statistical Sciences, 2(1):59–68, May 2010. CODEN ???? ISSN 0974-6811.

Ferreira:2012:ECM

[FFNM12] Sandra Saraiva Ferreira, Dário Ferreira, Célia Nunes, and João Tiago Mexia. Estimation and crossing for models with commutative orthogonal block structure. Advances and Applications in Statistical Sciences, 7(2):23–34, January 2012. CODEN ???? ISSN 0974-6811.

Figini:2010:BCM

[FG10] Silvia Figini and Paolo Giudici. Bayesian churn models. Advances and Applications in Statistical Sciences, 1(2):285–310, February 2010. CODEN ???? ISSN 0974-6811.

Fernandes:2013:SOR

[FLP13] Cristiano Fernandes, César Lima, and Adrian Pizzinga. A simulator for oil and refined products spot prices using a multivariate state space model. Advances and Applications in Statistical Sciences, 8(5):129–??, September 2013. CODEN ???? ISSN 0974-6811.

Fokoue:2012:BIS

[Fok12] Ernest Fokoué. Beta induced sparsity algorithm. Advances and Applications in Statistical Sciences, 7(7):75–82, October 2012. CODEN ???? ISSN 0974-6811.

Fotouhi:2010:JMA

[Fot10] Ali Reza Fotouhi. A joint model for analyzing longitudinal binary and event history data: a simulation study. Advances and Applications in Statistical Sciences, 4(1):61–71, November 2010. CODEN ???? ISSN 0974-6811.

Fotouhi:2011:EMD

[Fot11] Ali Reza Fotouhi. The effect of mixing distribution on the estimation of the parameters in random effects models. Advances and Applications in Statistical Sciences, 6(3):137–151, May 2011. CODEN ???? ISSN 0974-6811.

From:2010:CIB

[FS10a] Steven G. From and Andrew W. Swift. Convolution of independent Bernoulli random variables and some new approximations. Advances and Applications in Statistical Sciences, 2(1): 37–50, May 2010. CODEN ???? ISSN 0974-6811.

From:2010:GME

[FS10b] Steven G. From and Andrew W. Swift. Generalized median estimators for small even sample sizes. Advances and Applications in Statistical Sciences, 4(2):145–164, 2010. CODEN ???? ISSN 0974-6811.

Figliola:2011:FMS

[FS11] A. Figliola and E. Serrano. Fractal model for the stock market indices. Advances and Applications in Statistical Sciences, 6(6): 445–459, 2011. CODEN ???? ISSN 0974-6811.

Figliola:2010:MCS

[FZSR10] A. Figliola, L. Zunino, E. Serrano, and O. A. Rosso. Multifractal characterization of the stock market inefficiency using the wavelet leaders method. Advances and Applications in Statistical Sciences, 2(2):287–302, August 2010. CODEN ???? ISSN 0974-6811.

Ghahroodi:2013:CLC

[GG13] Z. Rezaei Ghahroodi and M. Ganjali. Consumption over the life cycle in Iran: application to the household expenditure and income survey during 1996–2008. Advances and Applications in Statistical Sciences, 8(2):49–66, 2013. CODEN ???? ISSN 0974-6811.

Ganjali:2010:WGE

[GGJ10] M. Ganjali, Z. Rezaei Ghahroodi, and M. Javadi. A weighted generalized estimating equations approach for analyzing clustered unemployment data. Advances and Applications in Statistical Sciences, 3(1):157–175, 2010. CODEN ???? ISSN 0974-6811.

Gioia:2011:PFM

[Gio11] Federica Gioia. Pricing in financial markets with interval data: arbitrage and order between intervals. Advances and Applications in Statistical Sciences, 6(6):597–613, September 2011. CODEN ???? ISSN 0974-6811.

Grira:2014:CLT

[Gri14] Sofiane Grira. Central limit theorems for martingales. Advances and Applications in Statistical Sciences, 9(1):1–16, January 2014. CODEN ???? ISSN 0974-6811.

Grira:2016:SCC

[Gri16] Sofiane Grira. On stable convergence in the central limit theorem. Advances and Applications in Statistical Sciences, 11(2): 9–17, April 2016. CODEN ???? ISSN 0974-6811.

Hossain:2010:GED

[HA10] Syed Afzal Hossain and Mohammad Ahsanullah. On generalized exponential distributions. Advances and Applications in Statistical Sciences, 4(1):1–22, November 2010. CODEN ???? ISSN 0974-6811.

Hall:2010:EEM

[HDS10] Owen P. Hall, Jr., Charles J. Delaney, and Samuel L. Seaman. Evaluating the efficacy of mandatory attendance. Advances and Applications in Statistical Sciences, 2(1):167–184, May 2010. CO-DEN ???? ISSN 0974-6811.

He:2012:OPF

[He12] Bingjie He. Optimal policy of fixed-period inventory control of HRMs based on Markov decision process. Advances and Applications in Statistical Sciences, 7(6):65–73, September 2012. CO-DEN ???? ISSN 0974-6811.

Hu:2010:IUD

[HGXA10] Xiaojun Hu, Gary L. Gadbury, Qinfang Xiang, and David B. Allison. Illustrations on using the distribution of a P -value in high dimensional data analyses. Advances and Applications in Statistical Sciences, 1(2):191–213, 2010. CODEN ???? ISSN 0974-6811.

Hurley:2011:MCS

[Hur11] W. J. Hurley. Managing the client-statistician interface: the case of admission decisions and elite varsity athletes at the Royal Military College of Canada. Advances and Applications in Statistical Sciences, 6(2):89–101, 2011. CODEN ???? ISSN 0974-6811.

Huson:2010:BBA

[Hus10] L. W. Huson. Bootstrapped Bayesian analyses of two adaptive clinical trials. Advances and Applications in Statistical Sciences, 4(1):23–30, 2010. CODEN ???? ISSN 0974-6811.

Hutson:2009:CPH

[Hut09] Alan D. Hutson. Calculation of the per hypothesis error rate via sums of Steck’s determinants. Advances and Applications in Statistical Sciences, 1(1):31–38, November 2009. CODEN ???? ISSN 0974-6811.

Heinrich:2010:SCE

[HW10] Lothar Heinrich and Michael Werner. Strongly consistent estimation in Wicksell’s corpuscle problem in case of occluded spheres. Advances and Applications in Statistical Sciences, 4(2):89–119, December 2010. CODEN ???? ISSN 0974-6811.

Ing:2010:ULP

[Ing10] Marsha Ing. Using latent profile analysis to model perceptions of school safety. Advances and Applications in Statistical Sciences, 3(2):237–253, 2010. CODEN ???? ISSN 0974-6811.

Jensen:2010:TME

[JR10] D. R. Jensen and D. E. Ramirez. Tracking MSE efficiencies in ridge regression. Advances and Applications in Statistical Sciences, 1(2):381–398, 2010. CODEN ???? ISSN 0974-6811.

Karabatsos:2009:MHS

[Kar09] George Karabatsos. Modeling heteroscedasticity in the singleindex model with the Dirichlet process. Advances and Applications in Statistical Sciences, 1(1):83–104, November 2009. CO-DEN ???? ISSN 0974-6811.

Kimiagar:2011:HPU

[Kim11] Salimeh Kimiagar. Hydrophilicity, photocatalytic and universality class of n-tio 2. Advances and Applications in Statistical Sciences, 6(5):329–??, August 2011. CODEN ???? ISSN 0974-6811.

Kuna:2010:PDU

[KMV10] Horacio Kuna, Ramón García Martínez, and Francisco R. Villatoro. Pattern discovery in university students desertion based on data mining. Advances and Applications in Statistical Sciences, 2(2):275–285, 2010. CODEN ???? ISSN 0974-6811.

Khan:2010:ARL

[KP10a] Hafiz M. R. Khan and Serge B. Provost. Application of the Rayleigh life model to a type II censored remission times data set. Advances and Applications in Statistical Sciences, 2(1):151– 165, 2010. CODEN ???? ISSN 0974-6811.

Khan:2010:BSI

[KP10b] Hafiz M. R. Khan and Serge B. Provost. Bayesian statistical inference for future responses from parallel systems. Advances and Applications in Statistical Sciences, 1(2):487–510, 2010. CODEN ???? ISSN 0974-6811.

Kozubowski:2010:NJD

[KP10c] Tomasz J. Kozubowski and Anna K. Panorska. A note on the joint distribution involving Poissonian sum of exponential variables.

Advances and Applications in Statistical Sciences, 1(2):157–165, 2010. CODEN ???? ISSN 0974-6811.

Kurita:2011:PEM

[Kur11] Takamitsu Kurita. A parsimonious econometric model of inflation-demand nexus in Japan. Advances and Applications in Statistical Sciences, 6(3):153–173, 2011. CODEN ???? ISSN 0974-6811.

Lefebvre:2010:ESK

[LA10] Mario Lefebvre and Djilali Ait Aoudia. Explicit solutions to the Kolmogorov backward equation. Advances and Applications in Statistical Sciences, 2(1):51–57, 2010. CODEN ???? ISSN 0974-6811.

Lee:2010:MEV

[LAB^{+}10] Kichun Sky Lee, M. Forrest Abouelnasr, Charlene W. Bayer, Sheryl G. A. Gabram, Boris Mizaikoff, André Rogatko, and Brani Vidakovic. Mining exhaled volatile organic compounds for breast cancer detection. Advances and Applications in Statistical Sciences, 1(2):327–342, February 2010. CODEN ???? ISSN 0974-6811.

Lanza:2010:TDU

[LB10] Stephanie T. Lanza and Bethany C. Bray. Transitions in drug use among high-risk women: an application of latent class and latent transition analysis. Advances and Applications in Statistical Sciences, 3(2):203–235, October 2010. CODEN ???? ISSN 0974-6811.

Lefebvre:2010:IS

[Lef10] Mario Lefebvre. Is the Internet slowing down? Advances and Applications in Statistical Sciences, 1(2):481–486, February 2010. CODEN ???? ISSN 0974-6811.

Lima:2010:IMD

[LLF10] F. W. S. Lima, Edina M. S. Luz, and R. N. Costa Filho. Ising model S = 1, 3/2 and 2 on directed networks. Advances and Applications in Statistical Sciences, 2(1):185–199, 2010. CODEN ???? ISSN 0974-6811.

LaCour:2011:TAF

[LM11a] Lisbeth Funding La Cour and Anders Milhøj. Temporal aggregation in first order cointegrated vector autoregressive models.

Advances and Applications in Statistical Sciences, 6(4):207–227, July 2011. CODEN ???? ISSN 0974-6811.

Levantesi:2011:MML

[LM11b] Susanna Levantesi and Massimiliano Menzietti. Modelling and managing longevity and disability risks in long term care insurance. Advances and Applications in Statistical Sciences, 6(6): 549–573, September 2011. CODEN ???? ISSN 0974-6811.

LaCour:2015:BTP

[LM15] Lisbeth La Cour and Anders Milhøj. Benefits from taking a private tutorial course for exam preparation? The case of macroeconomics at Copenhagen Business School. Advances and Applications in Statistical Sciences, 10(1):1–31, March 2015. CODEN ???? ISSN 0974-6811.

Lee:2010:URB

[LPPY10] Chu-In Charles Lee, Chul Gyu Park, Jianan Peng, and Xiaosong Yan. A unified representation of bivariate positive dependence. Advances and Applications in Statistical Sciences, 3(1):65–82, 2010. CODEN ???? ISSN 0974-6811.

LouzadaNeto:2010:HHM

[LTC10] Francisco Louzada Neto, Cynthia A. V. Tojeiro, and Gleici Castro Perdoná. A hybrid hazard model with threshold-stress: the Weibull case. Advances and Applications in Statistical Sciences, 1(2):145–155, February 2010. CODEN ???? ISSN 0974-6811.

Luo:2011:POM

[Luo11] June Luo. A penalized optimization method for classification. Advances and Applications in Statistical Sciences, 6(4):243–253, July 2011. CODEN ???? ISSN 0974-6811.

Luo:2012:SAF

[Luo12] June Luo. Statistical analysis for factors in relation to diabetes. Advances and Applications in Statistical Sciences, 7(7): 83–90, 2012. CODEN ???? ISSN 0974-6811.

Luo:2011:SMD

[LZ11] June Luo and Yijun Zuo. Statistical methods for differential gene detection in microarray. Advances and Applications in Statistical Sciences, 6(2):77–87, April 2011. CODEN ???? ISSN 0974-6811.

Mokaddis:2013:PAT

[MAAH13] G. S. Mokaddis, Y. M. Ayed, and H. S. Al-Hajeri. Probabilistic analysis of two warm standby system with instructions at need. Advances and Applications in Statistical Sciences, 8(3): 67–??, May 2013. CODEN ???? ISSN 0974-6811.

MacDougall:2011:MBN

[Mac11] Margaret Macdougall. Moving beyond the nuts and bolts of score reliability in medical education: some valuable lessons from measurement theory. Advances and Applications in Statistical Sciences, 6(7):643–664, 2011. CODEN ???? ISSN 0974-6811.

Maruyama:2009:NSM

[Mar09] Yosihito Maruyama. A note on sample measure of multivariate skewness. Advances and Applications in Statistical Sciences, 1(1): 39–44, 2009. CODEN ???? ISSN 0974-6811.

Marsh:2010:SGM

[Mar10a] Patrick Marsh. Some geometry for the maximal invariant in time series regressions. Advances and Applications in Statistical Sciences, 1(2):105–124, February 2010. CODEN ???? ISSN 0974-6811.

Martins:2010:MOC

[Mar10b] AndrÉ C. R. Martins. A middle option for choices in the continuous opinions and discrete actions model. Advances and Applications in Statistical Sciences, 2(2):333–345, 2010. CODEN ???? ISSN 0974-6811.

Maruyama:2010:MTC

[Mar10c] Yosihito Maruyama. On multivariate tests for comparing means. Advances and Applications in Statistical Sciences, 1(2):441–452, February 2010. CODEN ???? ISSN 0974-6811.

Massou:2011:IBC

[Mas11] Efthalia Massou. An information based criterion for analyzing supersaturated designs. Advances and Applications in Statistical Sciences, 6(1):59–69, 2011. CODEN ???? ISSN 0974-6811.

Mazzoni:2010:EPE

[Maz10] Thomas Mazzoni. Expected a posteriori estimation in finance. Advances and Applications in Statistical Sciences, 1(2):263–284, 2010. CODEN ???? ISSN 0974-6811.

Moreira:2010:MCC

[MCB10] Carla M. Moreira, Rosa M. Crujeiras, and Maria J. Bento. Mapping childhood cancer incidence at northern region of Portugal. Advances and Applications in Statistical Sciences, 3(1):177–193, September 2010. CODEN ???? ISSN 0974-6811.

Migueles:2010:MEP

[MLMB10] M. V. Migueles, C. Lagorio, P. A. Macri, and L. A. Braunstein. Modeling epidemics propagation on virtual cities. Advances and Applications in Statistical Sciences, 2(2):393–400, 2010. CODEN ???? ISSN 0974-6811.

Morlini:2010:NNA

[Mor10] Isabella Morlini. A nonlinear and nonparametric approach for ground level air pollutants forecasting. Advances and Applications in Statistical Sciences, 3(1):127–156, September 2010. CODEN ???? ISSN 0974-6811.

Moretto:2010:DEU

[MPT10] Enrico Moretto, Sara Pasquali, and Barbara Trivellato. Derivative evaluation using recombining trees under stochastic volatility. Advances and Applications in Statistical Sciences, 1(2):453–480, 2010. CODEN ???? ISSN 0974-6811.

Mohsenkhani:2011:IPP

[MR11] Zohreh Fallah Mohsenkhani and Mohammad Reza Farid Rouhani. Iran’s population projection using stochastic growth rate. Advances and Applications in Statistical Sciences, 5(1): 63–72, January 2011. CODEN ???? ISSN 0974-6811.

Maturo:2010:DMF

[MSV10] Antonio Maturo, Massimo Squillante, and Aldo G. S. Ventre. Decision making, fuzzy measures and hyperstructures. Advances and Applications in Statistical Sciences, 2(2):233–253, 2010. CODEN ???? ISSN 0974-6811.

Nandram:2011:HBM

[NBC11] Balgobin Nandram, Yan Bai, and Jai Won Choi. Hierarchical Bayesian models for assessing possible changes in prevalence of activity limitation. Advances and Applications in Statistical Sciences, 6(5):285–??, August 2011. CODEN ???? ISSN 0974-6811.

Najafi:2011:VAP

[NFZ11] H. Saberi Najafi, B. Fathi, and A. Zeinal. The variables affecting the performance of LDA. Advances and Applications in Statistical Sciences, 6(1):1–25, March 2011. CODEN ???? ISSN 0974-6811.

Nguyen:2011:EAT

[NG11] Ngoc Nguyen and Arjun K. Gupta. Estimation of allocative and technical efficiencies with fat-tailed error models. Advances and Applications in Statistical Sciences, 5(2):85–112, 2011. CODEN ???? ISSN 0974-6811.

Nylund-Gibson:2010:AML

[NGGJ10] Karen Nylund-Gibson, Sandra Graham, and Jaana Juvonen. An application of multilevel LCA to study peer victimization in middle school. Advances and Applications in Statistical Sciences, 3 (2):343–363, 2010. CODEN ???? ISSN 0974-6811.

Nagata:2010:CPA

[NN10] Koji Nagata and Tadao Nakamura. Can the Peano axioms meet Zermelo–Fraenkel set theory with the Axiom of Choice? Advances and Applications in Statistical Sciences, 3(1):195–201, 2010. CO-DEN ???? ISSN 0974-6811.

Olivera:2011:ISM

[OPA11] Noemi L. Olivera, Araceli N. Proto, and Marcel Ausloos. Information society: modeling a complex system with scarce data. Advances and Applications in Statistical Sciences, 6(6):505–??, September 2011. CODEN ???? ISSN 0974-6811.

Ouassou:2010:STE

[OR10] Idir Ouassou and Mustapha Rachdi. Stein type estimation of the regression operator for functional data. Advances and Applications in Statistical Sciences, 1(2):233–250, 2010. CODEN ???? ISSN 0974-6811.

Parisi:2010:BBP

[Par10] D. R. Parisi. Bulls and bears in a pedestrian counterflow system. Advances and Applications in Statistical Sciences, 2(2):347–357, August 2010. CODEN ???? ISSN 0974-6811.

Peng:2013:TES

[Pen13] Jianan Peng. Testing equivalence of several treatments with a control under simple tree ordering. Advances and Applications in Statistical Sciences, 8(6):163–??, October 2013. CODEN ???? ISSN 0974-6811.

Peugh:2010:GMM

[PF10] James Peugh and Xitao Fan. Growth mixture modeling: concepts and implementation. Advances and Applications in Statistical Sciences, 3(2):255–283, October 2010. CODEN ???? ISSN 0974-6811.

Perez-Lechuga:2011:MPS

[PLFRKT11] G. Pérez-Lechuga, C. F. Flores-Rivera, A. A. Karelyn, and A. Tarasenko. A multi period stochastic model to link the supply chain with the production plan in a manufacturing organization. Advances and Applications in Statistical Sciences, 6(4):255–283, 2011. CODEN ???? ISSN 0974-6811.

Proto:2010:FSI

[Pro10] Araceli N. Proto. Foreword [special issue: Proceedings of the IV Meeting on Dynamics of Social and Economic Systems]. Advances and Applications in Statistical Sciences, 2(2):i, 2010. CODEN ???? ISSN 0974-6811. Held in Pinamar, April 14–18, 2009, Special issue edited by Proto and A. Plastino.

Pasqualitto:2011:LRE

[PS11a] Emanuela Pasqualitto and Maria Sole Staffa. Longevity risk: an empirical analysis for life annuity. Advances and Applications in Statistical Sciences, 6(6):403–??, September 2011. CODEN ???? ISSN 0974-6811.

Proto:2011:FSI

[PS11b] Araceli N. Proto and Massimo Squillante. Foreword [Special issue: Proceedings of the V Meeting on Dynamics of Social and Economic Systems]. Advances and Applications in Statistical Sciences, 6(6):i, 2011. CODEN ???? ISSN 0974-6811. Held in Benevento, September 20–25, 2010.

Pang:2011:TEM

[PT11] Herbert Pang and Tiejun Tong. Testing equality of means when one sample has no replication. Advances and Applications in Statistical Sciences, 5(1):53–62, 2011. CODEN ???? ISSN 0974-6811.

Pizzinga:2011:DSA

[PVF11] Adrian Pizzinga, Luciano Vereda, and Cristiano Fernandes. A dynamic style analysis of exchange rate funds: the case of Brazil at the 2002 election. Advances and Applications in Statistical Sciences, 6(2):111–135, 2011. CODEN ???? ISSN 0974-6811.

Rotundo:2010:NAO

[RA10] G. Rotundo and A. M. D. Arcangelis. Network analysis of ownership and control structure in the Italian stock market. Advances and Applications in Statistical Sciences, 2(2):255–273, August 2010. CODEN ???? ISSN 0974-6811.

Raveh:2010:QRA

[RL10] Adi Raveh and Guy Leshem. A quick and rather accurate estimate of Pearson’s correlation coefficient ρ from scatter diagrams. Advances and Applications in Statistical Sciences, 1(2):371–380, February 2010. CODEN ???? ISSN 0974-6811.

Rancan:2010:IRB

[RPGM10] Claudio Rancan, Patricia Pesado, and Ramón García-Martínez. Issues in rule based knowledge discovering process. Advances and Applications in Statistical Sciences, 2(2):303–314, 2010. CODEN ???? ISSN 0974-6811.

Ranjbar:2011:BAP

[RRA11] Esmail Ranjbar, Vahid Ranjbar, and Mohammad Ghasem Akbarei. Bayesian approach to point estimation with fuzzy data. Advances and Applications in Statistical Sciences, 5(2):73–83, February 2011. CODEN ???? ISSN 0974-6811.

Soares:2010:CAR

[SC10] João Oliveira Soares and Cristina C. Coutinho. Cluster analysis in regional science. Advances and Applications in Statistical Sciences, 1(2):311–325, 2010. CODEN ???? ISSN 0974-6811.

Sellers:2012:DDD

[Sel12] Kimberly F. Sellers. A distribution describing differences in count data containing common dispersion levels. Advances and Applications in Statistical Sciences, 7(3):35–46, March 2012. CODEN ???? ISSN 0974-6811.

Shimshak:2010:IMP

[Shi10] Daniel G. Shimshak. Integration of managerial preferences in DEA modeling: a case study involving weight restrictions. Advances and Applications in Statistical Sciences, 1(2):167–190, February 2010. CODEN ???? ISSN 0974-6811.

Simonetti:2011:EPA

[SM11] Biagio Simonetti and Gabriella Marcarelli. Estimation of priorities in the AHP through taxicab decomposition. Advances and Applications in Statistical Sciences, 6(6):393–402, 2011. CODEN ???? ISSN 0974-6811.

Spencer:2015:MCS

[SM15] Bryan P. Spencer and Ganesh B. Malla. A Monte Carlo simulation comparison of some survival functions for incomplete data. Advances and Applications in Statistical Sciences, 10(2):33–43, May 2015. CODEN ???? ISSN 0974-6811.

Saraiva:2014:PAS

[SMLS14] Erlandson F. Saraiva, Luís A. Milan, Francisco Louzada, and Adriano K. Suzuki. A predictive allocation sampler for clustering of gene expression data. Advances and Applications in Statistical Sciences, 9(2):17–??, May 2014. CODEN ???? ISSN 0974-6811.

Sarris:2010:DMU

[SP10] C. M. Sarris and A. N. Proto. Decision making under stress: a quantum model approach. Advances and Applications in Statistical Sciences, 2(2):359–373, 2010. CODEN ???? ISSN 0974-6811.

Sarris:2011:DSN

[SP11] C. M. Sarris and A. N. Proto. Dissipative SU(2) semiquantum non-linear Hamiltonian for decision making. Advances and Applications in Statistical Sciences, 6(6):341–355, September 2011. CODEN ???? ISSN 0974-6811.

Spanos:2011:MCN

[Spa11a] Aris Spanos. Misplaced criticisms of Neyman–Pearson (N–P) testing in the case of two simple hypotheses. Advances and Applications in Statistical Sciences, 6(4):229–242, 2011. CODEN ???? ISSN 0974-6811.

Spanos:2011:RWU

[Spa11b] Aris Spanos. Revisiting the Welch uniform model: a case for conditional inference? Advances and Applications in Statistical Sciences, 5(1):33–52, January 2011. CODEN ???? ISSN 0974-

Spanos:2013:RME

[Spa13] Aris Spanos. Revisiting the ‘mixed experiment’ example: fallacious frequentist reasoning or an improper statistical model? Advances and Applications in Statistical Sciences, 8(2):29–47, January 2013. CODEN ???? ISSN 0974-6811.

Szybisz:2010:PCB

[SS10] Leszek Szybisz and Martín A. Szybisz. People’s collective behavior in a hyperinflation. Advances and Applications in Statistical Sciences, 2(2):315–331, August 2010. CODEN ???? ISSN 0974-6811.

Stoica:2010:RCP

[Sto10] George Stoica. Rate of convergence in Petrov’s strong law of large numbers. Advances and Applications in Statistical Sciences, 4(1): 57–60, 2010. CODEN ???? ISSN 0974-6811.

Tarsitano:2010:CST

[Tar10] Agostino Tarsitano. Classification of short time series. Advances and Applications in Statistical Sciences, 1(2):215–231, February 2010. CODEN ???? ISSN 0974-6811.

Tan:2010:LDA

[TBLW10] Qier Tan, Eric Bohm, Lynda Loucks, and Xikui Wang. Lifetime data analysis of waiting time to healthcare in Canada. Advances and Applications in Statistical Sciences, 2(1):89–100, May 2010. CODEN ???? ISSN 0974-6811.

Tolusso:2010:LLI

[TC10] David Tolusso and Richard J. Cook. Local likelihood for intervalcensored recurrent event data. Advances and Applications in Statistical Sciences, 1(2):343–369, 2010. CODEN ???? ISSN 0974-6811.

Thiagarajah:2010:QEE

[Thi10] K. Thiagarajah. Quantile estimation for the extreme value distribution with censored data. Advances and Applications in Statistical Sciences, 1(2):429–440, 2010. CODEN ???? ISSN 0974-6811.

Thiagarajah:2012:ARP

[Thi12] K. Thiagarajah. Approximate ruin probabilities when the claim size distribution is Weibull. Advances and Applications in Statistical Sciences, 7(5):55–64, July 2012. CODEN ???? ISSN 0974-6811.

Tang:2009:TET

[TT09] Nian-Sheng Tang and Man-Lai Tang. Testing the equality of two ordered multinomial proportions. Advances and Applications in Statistical Sciences, 1(1):57–82, 2009. CODEN ???? ISSN 0974-6811.

Tahata:2009:DSM

[TTT09] Kouji Tahata, Takanori Toyoda, and Sadao Tomizawa. The delta symmetry model for square contingency tables. Advances and Applications in Statistical Sciences, 1(1):25–30, 2009. CODEN ???? ISSN 0974-6811.

Tahata:2010:TMP

[TTT10] Kouji Tahata, Hiroki Tokuno, and Sadao Tomizawa. On test of marginal point-symmetry in multi-way tables. Advances and Applications in Statistical Sciences, 3(1):1–12, September 2010. CODEN ???? ISSN 0974-6811.

Tong:2010:CCH

[TWW10] Q. Tong, X. Wang, and Y. H. Wu. Combining current and historical survival data using weighted likelihood. Advances and Applications in Statistical Sciences, 2(1):129–149, May 2010. CODEN ???? ISSN 0974-6811.

Varron:2010:NUF

[Var10] Davit Varron. A nonstandard uniform functional limit law for the increments of the multivariate empirical distribution function. Advances and Applications in Statistical Sciences, 1(2):399–428, February 2010. CODEN ???? ISSN 0974-6811.

Vasilescu:2010:MRI

[Vas10] Cezar Vasilescu. Modeling the reliability of information management systems based on mission specific tools set software. Advances and Applications in Statistical Sciences, 4(2):121–134, 2010. CODEN ???? ISSN 0974-6811.

Walls:2010:GSM

[Wal10] W. D. Walls. General stable models of the rate of return to Hollywood films. Advances and Applications in Statistical Sciences, 2(1):19–36, 2010. CODEN ???? ISSN 0974-6811.

Wang:2009:GGT

[Wan09] Jichuan Wang. Gauging growth trajectories of two related outcomes simultaneously: an application of parallel latent growth modeling. Advances and Applications in Statistical Sciences, 1 (1):1–23, November 2009. CODEN ???? ISSN 0974-6811.

Wang:2010:LSG

[Wan10] Jichuan Wang. Longitudinal study on growth trajectory of crackcocaine use: applications of group-based trajectory models. Advances and Applications in Statistical Sciences, 3(2):285–310, 2010. CODEN ???? ISSN 0974-6811.

Wang:2011:GDR

[Wan11] Jiantian Wang. Generalized t-distributions from regressional property. Advances and Applications in Statistical Sciences, 6 (2):103–109, April 2011. CODEN ???? ISSN 0974-6811.

Wang:2010:PSI

[WL10] Jichuan Wang and Stephanie T. Lanza. Preface [special issue: Methods and applications of mixture models]. Advances and Applications in Statistical Sciences, 3(2):1–6, 2010. CODEN ???? ISSN 0974-6811.

Willie:2010:DMH

[WN10a] Michael Mncedisi Willie and Phakamile Nkomo. Dealing with multicollinearity in healthcare benefit options. Advances and Applications in Statistical Sciences, 3(1):113–125, 2010. CODEN ???? ISSN 0974-6811.

Willie:2010:LMM

[WN10b] Mncedisi Michael Willie and Phakamile Nkomo. Linear mixed models: an application to medical schemes contributions data. Advances and Applications in Statistical Sciences, 2(1):201–211, May 2010. CODEN ???? ISSN 0974-6811.

Webb:2010:IUH

[WSNP10] Mi-Young L. Webb, Paula J. Schwanenflugel, and Stacey Neuharth-Pritchett. Investigating unobserved heterogeneity in teacher-child relationships: an application of latent class factor analysis. Advances and Applications in Statistical Sciences, 3(2): 311–341, October 2010. CODEN ???? ISSN 0974-6811.

Wu:2011:EBD

[WW11] Yuehua Wu and Xiaolan L. Wang. M-estimation based detection of undocumented mean shifts without trend changes for the Box– Cox transformed daily precipitation data series. Advances and Applications in Statistical Sciences, 6(3):187–205, 2011. CODEN ???? ISSN 0974-6811.

Wu:2010:GNA

[WXZ10] Xin Wu, Miao Xia, and Peibiao Zhao. The generalized noarbitrage analysis of frictional markets. Advances and Applications in Statistical Sciences, 1(2):251–262, February 2010. CO-DEN ???? ISSN 0974-6811.

Xie:2011:FPT

[XK11] Shengkun Xie and Sridhar Krishnan. First passage times density of Markov switching geometric Brownian motion. Advances and Applications in Statistical Sciences, 5(1):1–22, January 2011. CODEN ???? ISSN 0974-6811.

Yin:2011:AIF

[Yin11] Xiangfeng Yin. An anticipating Itô formula for general Lévy processes with forward integral. Advances and Applications in Statistical Sciences, 5(1):23–31, 2011. CODEN ???? ISSN 0974-6811.

Yamamoto:2012:MMH

[YKT12] Kouji Yamamoto, Takenori Kimbara, and Sadao Tomizawa. Measure for marginal homogeneity based on distance powerdivergence in square contingency tables. Advances and Applications in Statistical Sciences, 7(4):47–54, May 2012. CODEN ???? ISSN 0974-6811.

Ye:2010:NAS

[YW10] Daxiang Ye and Qunying Wu. A note on the almost sure central limit theorem. Advances and Applications in Statistical Sciences, 4(1):79–88, November 2010. CODEN ???? ISSN 0974-6811.

Zabaleta:2010:QSD

[ZA10] O. G. Zabaleta and C. M. Arizmendi. Quantum search for the dating market. Advances and Applications in Statistical Sciences, 2(2):401–407, August 2010. CODEN ???? ISSN 0974-6811.

Zabaleta:2011:QDT

[ZA11] O. G. Zabaleta and C. M. Arizmendi. Quantum decision theory on a dating market. Advances and Applications in Statistical Sciences, 6(6):489–496, September 2011. CODEN ???? ISSN 0974-6811.

Zhao:2011:NWC

[ZRQ11] Yan D. Zhao, Dewi Rahardja, and Yongming Qu. A note on weak consistency of the elementwise weighted total least squares estimator in a multivariate measurement error model under mild conditions. Advances and Applications in Statistical Sciences, 6 (1):71–76, March 2011. CODEN ???? ISSN 0974-6811.

Zhang:2010:SBE

[ZZ10] Jiajia Zhang and Hongmei Zhang. A semiparametric Bayesian estimation method of the accelerated failure time model. Advances and Applications in Statistical Sciences, 1(2):125–144, 2010. CO-DEN ???? ISSN 0974-6811.